OpenStudy (anonymous):

Solve for n. k = 6n - 5 ---------- 4 A.6k-5 -------- 4 = N B. 4k +5 __________ 6 =N C.4k -5 --------- 6 =N D.6k+5 --------- 4 =N

OpenStudy (anonymous):

B

OpenStudy (kropot72):

k=(6n-5)/4 Multiplying both sides by 4 gives: 4k=6n-5 Adding 5 to both sides gives: 4k+5=6n-5+5=6n Dividing both sides by 6 gives: (4k+5)/6=n